OpenStudy (anonymous):

simplified form 4y^2-1 divided by 2y^2-3y-2 I think the answer is 2y-1divided by y-2 what would be the restriction?

OpenStudy (anonymous):

(4y^(2)-1)/((2y^(2)-3y-2) (4y^(2)-1) = (2y-1)(2y+1) (2y^(2)-3y-2) = (y - 2)(2y + 1) (2y-1)(2y+1)/(y - 2)(2y + 1) = (2y-1)/(y-2) looks about right but you should try division  Thus the simplified answer is 3/(y-2) + 2